IMPROVEMENT OF DATA PROCESSING METHODS IN MODERN FT-IR SPECTROSCOPY

\underline{SERGEY V. SHILOV}, Bruker Optics, 19 Fortune Dr., Billerica, MA 01821 USA; A. KEENS, A. SI-MON, Bruker Optik GmbH, D-76275 Ettlingen, Germany.

Advances in digital signal processing, and progress in Audio-CD and DVD recording, have brought ”Delta-Sigma” ADCs with higher resolutions, but with the apparent disadvantage of a fixed digitization rate. These new ADCs were utilized to implement additional functionality, such as compensation of the detector’s frequency response and compensation of source fluctuations. Further advantages of ”Delta-Sigma” ADCs will be discussed with reference to signal-to-noise ratio for modern FT-IR spectrometers.